# ROTATING NEUTRON STARS WITH CHAOTIC MAGNETIC FIELDS IN GENERAL RELATIVITY AND RASTALL GRAVITY

### v9

@@ 1 Introduction
It has been widely understood that the evolution of the massive stars whose mass is 8--25 $M_{\odot}$ ends when neutron stars (NSs) form. The corresponding formation is due to the stars’ gravitational collapse during Type-II, Ib, or Ic supernova explosion phenomena [1, 2]. NSs are compact remnants of evolved stars, with degenerate fermions compressed by strong gravitational fields [3]. They are compact enough to provide necessary conditions for exotic physics to occur [4]. The stabilizing effect of gravity allows long-time-scale weak interactions (such as electron captures) to reach equilibrium, forming matter that is neutron-rich [5]. These stars cover an extensive density range, roughly several times of nuclear density ($n_{0}\sim 10^{14}$ gm/cc), throughout their outer crust to their inner core [6]. NSs are normally rotating as all celestial objects tend to naturally always rotate due to the law of angular momentum conservation [7]. When rotating NSs are highly magnetized, they are recognized as magnetars [8, 9, 10, 11, 12].
NSs have a wide range of magnetic fields. Observations indicate that magnetic fields on NSs are at least in the range of $10^{8-15}$ G [3]. It is important to note that the direct observation of the magnetic field in NSs’ cores remains unachieved. However, theoretical prediction suggests that the strengths of the magnetic fields in NSs’ cores are in the order of $10^{18}$ G to $10^{20}$ G [13, 14]. New classes of pulsars such as Anomalous X-ray pulsars (AXPs) and Soft-Gamma Repeater (SGRs) have been identified producing vast magnetic fields. SGR is associated with remnants of a supernova, which is a young NS [15].
@@
The works that has been mention above are constructed upon general relativity (GR). Moreover, the works presented in Refs. [28, 29, 30], which deal with rotating NSs with magnetic fields, are also based on GR. However, if GR is correct, there are some problems left to be explained, e.g. dark matter and dark energy [31, 32]. Dark energy cannot be consistently described solely by GR. The reason is that the dark energy EOS $\omega_{DE}$ is constrained to have values $\omega_{DE}=-1.018\pm 0.031$. This fact means that it is a probability that dark energy is actually phantom dark energy. In the context of GR, phantom dark energy is described only by phantom scalars, which do not provide a physically acceptable description of nature [33]. In addition, the Hubble tension—referring to the discrepancy in the measurements of the rate of the expansion of the universe, when comparing local measurements with cosmic microwave background radiation observations —remains another unresolved issue within the framework of GR [34]. There are also some problems related to compact objects. For example, the problem of the maximum allowed mass of horizonless compact objects exceeded GR predictions [35]. As the consequence, one might consider modified theories of gravity as the altenative ways. One of them is Rastall gravity (RG) [36].
@@ 2.1 Nuclear Model of the EOS
All mathematical formulations regarding the EOS used in this work is based on Ref. [68]. The QHD Lagrangian is given by
@@
$(g_{N\sigma}/m_{\sigma})^{2}$ 11.785 fm2 $n_{0}$ (fm-3) 0.1480.1700.148-0.1700.148 - 0.170 0.153 $M_{\Lambda}=1116$
$(g_{N\omega}/m_{\omega})^{2}$ 7.148 fm2 $M^{*}/M$ 0.70.80.7-0.80.7 - 0.8 0.7 $M_{\Sigma}=1193$
$(g_{N\rho}/m_{\rho})^{2}$ 3.880 fm2 $K$ (--MeV) 200315200-315200 - 315 300 $M_{\Xi}=1318$
$\kappa/M_{N}$ 0.0058940.0058940.005894 $S_{0}$ (MeV) 303430-3430 - 34 30.5 $m_{e}=0.511$
$\lambda$ 0.006426-0.006426- 0.006426 $B/A$ (MeV) 15.716.515.7-16.515.7 - 16.5 16.3 $m_{\mu}=105.6$
$M_{N}$ 939939939 MeV $L$ (MeV) 3611336-11336 - 113 88 --
@@
$\left[\gamma_{0}(i\partial^{0}-g_{b,\omega}\omega_{0}-g_{b,\rho}\rho_{0})-\gamma_{j}(i\partial^{j}-e_{b}A^{j})-M_{b}^{*}\right]\Psi$ $=$ 0 0,0 , (8)
@@ 2.2 Chaotic Magnetic Field of Neutron Stars
Figure 1: Relation between total energy density $\varepsilon$ and total pressure $p$.
@@ 2.4 Hartle-Thorne Formalism for Rotating Relativistic Stars
[…]
@@
Now we proceed to the quadrupole sector which is not considered in Ref. [1]. From the ($\theta\theta$), ($\phi$,$\phi$), ($\theta r$), and ($rr$) components of the $l=2$ order of Einstein field equation, we obtain
@@
×$\times\frac{dj^{2}}{dr}-\frac{2}{3}\frac{\kappa\lambda}{4\kappa\lambda-1}\left(\frac{dj^{2}}{dr}\right)r\bar{\omega}^{2}\left[\frac{r}{(r-2m)\left(2\frac{d\varphi}{dr}\right)}\right]\frac{d\varepsilon}{dp}.$
@@ 3 Results and Discussion
We numerically solved all differential equations using Euler method. For the value of $\Omega$, we consider $\Omega=1000$ s-1 and $\Omega=2000$ s-1. However, except for the calculation of the moment of inertia, we use $\Omega=50$ s-1, as we only employ the very slow rotation approximation for this purpose. For the initial value of $\bar{\omega}$ (i.e., $\omega_{c}$), we adopt $\omega_{c}=80$ s-1 in all cases throughout this work. It is worth noting that the value $\omega_{c}=80$ used in our calculations is significantly smaller than the value used in Ref. [84], where $\bar{\omega}=3000$ s-1. Moreover, it is also worth noting that $\bar{\omega}(r)$ increases with $r$, which would result in a large $\bar{\omega}$ value at the surface if a large initial $\omega_{c}$ is chosen. On the other side, we have insight that according to Ref. [85], the HT formalism is accurate to better than 1 per cent even for the fastest millisecond pulsars. However, putting $\bar{\omega}_{c}$ = 3000 s-1 in the calculation is still potential to be dangerous, for the HT formalism is basically an approximation for slowly rotating relativistic stars. So our value of $\bar{\omega}_{c}$ is safer than the one in Ref. [84].
@@
As shown in Appendix Appendix: Macroscopic properties of neutron stars under different configurations., variations in $\gamma$ affect the $\langle R\rangle$ at both the maximum mass and the canonical mass only at the order of $10^{-3}$, which represents the largest magnitude of influence found—and even then, only for a subset of the configurations. Most of the values remain unchanged. In fact, the maximum mass is entirely unaffected by variations in $\gamma$, except for the configuration with $\lambda=-0.6\times 10^{-5}$ and $\Omega=0$, where the maximum mass is 1.99 $M_{\odot}$ for $\gamma=3$ and 1.98 $M_{\odot}$ for $\gamma=5$. This indicates that the impact of $\gamma$ on the overall stellar structure is minimal within the explored parameter range. Similar results for the same parameter range have also been reported in Refs. [24, 68, 87]. The appearance of the magnetic field is generally decrease the eccentricity, which means that generally, the deformation is weakened once the magnetic field appears.
In the context of Rastall parameter, the greater values of $\lambda$ results in more compact NSs as the radii decrease, in which this fact is obviously presented in Fig. 2. The greater values of Rastall parameter also cause the eccentricity diminished, in which this result can be validated in Fig. 5. In the term of the rotation, as we expect, faster rotations result in the increment of the eccentricity, since the deformation becomes stronger as the rotation getting faster.
Fig. 2 presents the mass-radius relation of NSs in a static configuration. This relation purely comes from TOV equation. From the relation, we can see that the presence of a chaotic magnetic field can increase the maximum mass of the NSs and simultaneously reduce the radius of NSs, as indicated by the leftward shift of the curves with the appearance of magnetic field strength, suggesting that a chaotic magnetic field can contribute to greater compactness of the stars. Furthermore, greater values of Rastall parameter lead to the decrease of the radii of NSs which also directly lead to the increment of NSs’ compactness. Compactness, defined as the mass-to-radius ratio, is directly influenced by this field. Moreover, the appearece of the chaotic magnetic field results in the mass of NSs reach the mass range of J0348+0432, i.e. $2.01\pm 0.04\>M_{\odot}$ [88], shown by yellow rectangle. All mass curve also in the agreement with the radius range of the canonical mass 1.4 $M_{\odot}$, shown by light blue line. The radius range of the canonical mass is taken from Ref. [82], which is based on the observational results reported in Refs. [89, 90], i.e. 11.5211.5211.52 km $<R_{1.4}<$ 14.2614.2614.26 km.
@@
We can see that at lower masses, the curves of moment of inertia tend to coincide, suggesting that the impact of chaotic magnetic field strength on the moment of inertia is negligible in this regime. However, a noticeable separation appears within the mass range $M=1.5-1.99$ $M_{\odot}$, where the appearance of chaotic magnetic fields lead to a increment in the moment of inertia. On the other hand, we also can see Rastall parameter has little impact on the moment of inertia of the NSs.
Remarkably, our findings for the moment of inertia are consistent with the constraint range provided by Landry et al. [27]. It is worth noting that this constraint evaluates whether rotating NSs with chaotic magnetic fields align with observational evidence. The constraint is given as $I=1.78^{+0.66}_{-0.66}\times 10^{45}$ g cm2 at $M=1.4$ $M_{\odot}$, the standard canonical mass of NSs. As detailed in Ref. [27], this constraint is derived from adio observations of massive pulsars.
@@ 4 Conclusion
In this study, we extend the formulation of rotating NSs with chaotic magnetic field in RG theory, by considering the effect of Rastall parameter in the quadrupole sector of HT formalism. We calculated the mass and radius of NSs under the influence of a chaotic magnetic field within GR and RG framework, enabling us to obtain the mass-radius relationship for these stars. The magnetic field ansatz used in this work is the one that is coupled to the energy density, which is propose in Ref. [24]. This ansatz behaves as parameter-free when $\gamma\geq 2$. We use three configurations: (i) $B=0$; (ii) $B_{0}=3\times 10^{18}$ G and $B_{\text{surf}}=10^{15}$ G with $\gamma=3$; and (iii) $B_{0}=3\times 10^{18}$ G and $B_{\text{surf}}=10^{15}$ G with $\gamma=5$. In the context of RG, we use 5 values of Rastall parameter $\lambda$, i.e. $\lambda=-1\times 10^{-5}$, $\lambda=-0.6\times 10^{-5}$, $\lambda=0$ (GR), $\lambda=0.6\times 10^{-5}$, and $\lambda=1\times 10^{-5}$. For the rotation, we employ three values of $\Omega$, i.e. $\Omega=0$ (static configuration), $\Omega=1000$ s-1, and $\Omega=2000$ s-1.
Our findings indicate that both magnetic fields and greater values of Rastall parameter can decrease the radius of NSs. NSs formed with chaotic magnetic fields exhibit a higher maximum mass compared to those without chaotic magnetic fields. Additionally, the chaotic magnetic fields and greater values of Rastall parameter can contribute to a reduction in radii. This fact shows that both the chaotic magnetic fields and larger values of Rastall parameter can simultaneously enhance the compactness. Moreover, both factors weaken the deformation of NSs. Thus, chaotic magnetic fields and greater values of Rastall parameter decrease the values of eccentricity of the stars. In the term of angular velocity $\Omega$, higher value of $\Omega$ strengthens the stellar deformation, as we expect. For deformed rotating NSs, significant deformations occur at lower mass values.
@@ Acknowledgments
We sincerely thank Prof. Luiz L. Lopes for sharing the EOS data. We also acknowledge Tsuyoshi Miyatsu and Ryan Rizaldy for valuable discussions. We are grateful to Anna Campoy Ordaz for making her code publicly available, which allowed MLP to modify it for use in this work. MLP acknowledges the Indonesia Endowment Fund for Education (LPDP) for financial support. FPZ and GH would like to thank Kemendiktisaintek (The Ministry of Higher Education, Science, and Technology) of the Republic of Indonesia (through LPPM ITB, Indonesia) for partially financial support. HLP and MLP would like to thank the members of the Theoretical Physics Groups at Institut Teknologi Bandung for their hospitality. MFARS is supported by the Second Century Fund (C2F), Chulalongkorn University, Thailand.

### v10

@@ 1 Introduction
It has been widely understood that the evolution of the massive stars whose mass is 8-25 $M_{\odot}$ ends when neutron stars (NSs) form. The corresponding formation is due to the stars’ gravitational collapse during Type-II, Ib, or Ic supernova explosion phenomena [1, 2]. NSs are compact remnants of evolved stars, with degenerate fermions compressed by strong gravitational fields [3]. They are compact enough to provide necessary conditions for exotic physics to occur [4]. The stabilizing effect of gravity allows long-time-scale weak interactions (such as electron captures) to reach equilibrium, forming matter that is neutron-rich [5]. These stars cover an extensive density range, roughly several times of nuclear density ($n_{0}\sim 10^{14}$ gm/cc), throughout their outer crust to their inner core [6]. NSs are normally rotating as all celestial objects tend to naturally always rotate due to the law of angular momentum conservation [7]. When rotating NSs are highly magnetized, they are recognized as magnetars [8, 9, 10, 11, 12].
NSs have a wide range of magnetic fields. Observations indicate that magnetic fields on NSs are at least in the range of $10^{8-15}$ G [3]. It is important to note that the direct observation of the magnetic field in NSs’ cores remains unachieved. However, theoretical prediction suggests that the strengths of the magnetic fields in NSs’ cores are in the order of $10^{18}$ G to $10^{20}$ G [13, 14]. New classes of pulsars such as Anomalous X-ray pulsars (AXPs) and Soft-Gamma Repeaters (SGRs) have been identified producing vast magnetic fields. SGR is associated with remnants of a supernova, which is a young NS [15].
@@
The works that has been mention above are constructed upon general relativity (GR). Moreover, the works presented in Refs. [28, 29, 30], which deal with rotating NSs with magnetic fields, are also based on GR. However, if GR is correct, there are some problems left to be explained, e.g. dark matter and dark energy [31, 32]. Dark energy cannot be consistently described solely by GR. The reason is that the dark energy EOS $\omega_{DE}$ is constrained to have values $\omega_{DE}=-1.018\pm 0.031$. This fact means that it is a probability that dark energy is actually phantom dark energy. In the context of GR, phantom dark energy is described only by phantom scalars, which do not provide a physically acceptable description of nature [33]. In addition, the Hubble tension—referring to the discrepancy in the measurements of the rate of the expansion of the universe, when comparing local measurements with cosmic microwave background radiation observations —remains another unresolved issue within the framework of GR [34]. There are also some problems related to compact objects. For example, the problem of the maximum allowed mass of horizonless compact objects exceeded GR predictions [35]. As the consequence, one might consider modified theories of gravity as the alternative ways. One of them is Rastall gravity (RG) [36].
@@
Especially in theoretical astrophysics, several studies have reported notable findings by incorporating RG. Pattersons et al. [49] showed that although RG only slightly increases the mass of anisotropic NSs, the resulting stellar mass can still reach the upper bound of the mass range for the secondary compact object of GW190814 through the contribution of the Rastall parameter. It is important to note that the secondary component of GW190814 [66] has been hypothesized to be the most massive NS ever observed [67]. Consistently, da Silva et al. [25] found that NSs in RG can attain this upper limit, while Banerjee et al. [68] demonstrated that quark stars within RG may also reach it.
@@ 2.1 Nuclear Model of the EOS
All mathematical formulations regarding the EOS used in this work are based on Ref. [71]. The QHD Lagrangian is given by
@@
$(g_{N\sigma}/m_{\sigma})^{2}$ 11.785 fm2 $n_{0}$ (fm-3) 0.1480.1700.148-0.170 0.153 $M_{\Lambda}=1116$
$(g_{N\omega}/m_{\omega})^{2}$ 7.148 fm2 $M^{*}/M$ 0.70.80.7-0.8 0.7 $M_{\Sigma}=1193$
$(g_{N\rho}/m_{\rho})^{2}$ 3.880 fm2 $K$ (-MeV) 200315200-315 300 $M_{\Xi}=1318$
$\kappa/M_{N}$ $0.005894$ $S_{0}$ (MeV) 303430-34 30.5 $m_{e}=0.511$
$\lambda$ 0.006426-0.006426 $B/A$ (MeV) 15.716.515.7-16.5 16.3 $m_{\mu}=105.6$
$M_{N}$ $939$ MeV $L$ (MeV) 3611336-113 88 -
@@
$\left[\gamma_{0}(i\partial^{0}-g_{b,\omega}\omega_{0}-g_{b,\rho}\rho_{0})-\gamma_{j}(i\partial^{j}-e_{b}A^{j})-M_{b}^{*}\right]\Psi$ $=$ $0,$ (8)
@@ 2.2 Chaotic Magnetic Field of Neutron Stars
Figure 1: Relation between total energy density $\varepsilon$ and total pressure pp.
@@ 2.4 Hartle-Thorne Formalism for Rotating Relativistic Stars
$L=\frac{\kappa}{3}\int^{R}_{0}dr\>r^{4}\frac{\varepsilon+p}{\left(1-\frac{2m}{r}\right)^{1/2}}\>\bar{\omega}\>e^{-\varphi}.$ (48)
@@
Now we proceed to the quadrupole sector which is not considered in Ref. [1]. From the ($\theta\theta$), ($\phi$$\phi$), ($\theta r$), and ($rr$) components of the $l=2$ order of Einstein field equation, we obtain
@@
$\times\frac{dj^{2}}{dr}-\frac{2}{3}\frac{\kappa\lambda}{4\kappa\lambda-1}\left(\frac{dj^{2}}{dr}\right)r\bar{\omega}^{2}\left[\frac{r}{(r-2m)\left(2\frac{d\varphi}{dr}\right)}\right]\frac{d\varepsilon}{dp}.$
@@ 3 Results and Discussion
We numerically solved all differential equations using Euler method. For the value of $\Omega$, we consider $\Omega=1000$ s-1 and $\Omega=2000$ s-1. However, except for the calculation of the moment of inertia, we use $\Omega=50$ s-1, as we only employ the very slow rotation approximation for this purpose. For the initial value of $\bar{\omega}$ (i.e., $\omega_{c}$), we adopt $\omega_{c}=80$ s-1 in all cases throughout this work, except for the calculation of the moment of inertia, where $\omega_{c}=0$. It is worth noting that the value $\omega_{c}=80$ used in our calculations is significantly smaller than the value used in Ref. [87], where $\bar{\omega}=3000$ s-1. Moreover, it is also worth noting that $\bar{\omega}(r)$ increases with $r$, which would result in a large $\bar{\omega}$ value at the surface if a large initial $\omega_{c}$ is chosen. On the other side, we have insight that according to Ref. [88], the HT formalism is accurate to better than 1 per cent even for the fastest millisecond pulsars. However, putting $\bar{\omega}_{c}$ = 3000 s-1 in the calculation is still potential to be dangerous, for the HT formalism is basically an approximation for slowly rotating relativistic stars. So our value of $\bar{\omega}_{c}$ is safer than the one in Ref. [87].
@@
As shown in Appendix Appendix: Macroscopic properties of neutron stars under different configurations., variations in $\gamma$ affect the $\langle R\rangle$ at both the maximum mass and the canonical mass only at the order of $10^{-3}$, which represents the largest magnitude of influence found—and even then, only for a subset of the configurations. Most of the values remain unchanged. In fact, the maximum mass is entirely unaffected by variations in $\gamma$, except for the configuration with $\lambda=-0.6\times 10^{-5}$ and $\Omega=0$, where the maximum mass is 1.99 $M_{\odot}$ for $\gamma=3$ and 1.98 $M_{\odot}$ for $\gamma=5$. This indicates that the impact of $\gamma$ on the overall stellar structure is minimal within the explored parameter range. Similar results for the same parameter range have also been reported in Refs. [24, 71, 90]. The appearance of the magnetic field generally decreases the eccentricity, which means that generally, the deformation is weakened once the magnetic field appears.
In the context of Rastall parameter, the greater values of $\lambda$ result in more compact NSs as the radii decrease, in which this fact is obviously presented in Fig. 2. The greater values of Rastall parameter also cause the eccentricity diminished, in which this result can be validated in Fig. 5. It is important to note that the Rastall parameter exerts only a negligible influence on the maximum mass of NSs. A comparable behavior, where modified gravity has little effect on the stellar mass—particularly on the maximum mass—has also been reported in Ref. [91] (see Figs. 4 and 5 therein), where the NS mass is expressed as the ADM mass within the modified gravity framework adopted in their analysis.
In term of the rotation, as we expect, faster rotations result in the increment of the eccentricity, since the deformation becomes stronger as the rotation is getting faster.
Fig. 2 presents the mass-radius relation of NSs in a static configuration. This relation purely comes from TOV equation. From the relation, we can see that the presence of a chaotic magnetic field can increase the maximum mass of the NSs and simultaneously reduce the radius of NSs, as indicated by the leftward shift of the curves with the appearance of magnetic field strength, suggesting that a chaotic magnetic field can contribute to greater compactness of the stars. Furthermore, greater values of Rastall parameter lead to the decrease of the radii of NSs which also directly lead to the increment of NSs’ compactness. Compactness, defined as the mass-to-radius ratio, is directly influenced by this field. Moreover, the appearece of the chaotic magnetic field results in the mass of NSs reach the mass range of J0348+0432, i.e. $2.01\pm 0.04\>M_{\odot}$ [92], shown by yellow rectangle. All mass curves are also in the agreement with the radius range of the canonical mass 1.4 $M_{\odot}$, shown by light blue line. The radius range of the canonical mass is taken from Ref. [85], which is based on the observational results reported in Refs. [93, 94], i.e. $11.52$ km $R_{1.4}<$ $14.26$ km.
@@
We can see that at lower masses, the curves of moment of inertia tend to coincide, suggesting that the impact of chaotic magnetic field strength on the moment of inertia is negligible in this regime. However, a noticeable separation appears within the mass range M=1.51.99M=1.5-1.99 $M_{\odot}$, where the appearance of chaotic magnetic fields lead to an increment in the moment of inertia. On the other hand, we also can see Rastall parameter has little impact on the moment of inertia of the NSs.
Remarkably, our findings for the moment of inertia are consistent with the constraint range provided by Landry et al. [27]. It is worth noting that this constraint evaluates whether rotating NSs with chaotic magnetic fields align with observational evidence. The constraint is given as $I=1.78^{+0.66}_{-0.66}\times 10^{45}$ g cm2 at $M=1.4$ $M_{\odot}$, the standard canonical mass of NSs. As detailed in Ref. [27], this constraint is derived from radio observations of massive pulsars.
@@ 4 Conclusion
In this study, we extend the formulation of rotating NSs with chaotic magnetic field in RG theory, by considering the effect of Rastall parameter in the quadrupole sector of HT formalism. We calculated the mass and radius of NSs under the influence of a chaotic magnetic field within GR and RG framework, enabling us to obtain the mass-radius relationship for these stars. The magnetic field ansatz used in this work is the one that is coupled to the energy density, which is proposed in Ref. [24]. This ansatz behaves as parameter-free when $\gamma\geq 2$. We use three configurations: (i) $B=0$; (ii) $B_{0}=3\times 10^{18}$ G and $B_{\text{surf}}=10^{15}$ G with $\gamma=3$; and (iii) $B_{0}=3\times 10^{18}$ G and $B_{\text{surf}}=10^{15}$ G with $\gamma=5$. In the context of RG, we use 5 values of Rastall parameter $\lambda$, i.e. $\lambda=-1\times 10^{-5}$, $\lambda=-0.6\times 10^{-5}$, $\lambda=0$ (GR), $\lambda=0.6\times 10^{-5}$, and $\lambda=1\times 10^{-5}$. For the rotation, we employ three values of $\Omega$, i.e. $\Omega=0$ (static configuration), $\Omega=1000$ s-1, and $\Omega=2000$ s-1.
Our findings indicate that both magnetic fields and greater values of Rastall parameter can decrease the radius of NSs. NSs formed with chaotic magnetic fields exhibit a higher maximum mass compared to those without chaotic magnetic fields. Additionally, the chaotic magnetic fields and greater values of Rastall parameter can contribute to a reduction in radii. This fact shows that both the chaotic magnetic fields and larger values of Rastall parameter can simultaneously enhance the compactness. Moreover, both factors weaken the deformation of NSs. Thus, chaotic magnetic fields and greater values of Rastall parameter decrease the values of eccentricity of the stars. In term of angular velocity $\Omega$, higher value of $\Omega$ strengthens the stellar deformation, as we expect. For deformed rotating NSs, significant deformations occur at lower mass values.
@@ Acknowledgments
We sincerely thank Prof. Luiz L. Lopes for sharing the EOS data. We also acknowledge Tsuyoshi Miyatsu and Ryan Rizaldy for valuable discussions. We are grateful to Anna Campoy Ordaz for making her code publicly available, which allowed MLP to modify it for use in this work. MLP acknowledges the Indonesia Endowment Fund for Education (LPDP) for financial support. FPZ and GH would like to thank Kemendiktisaintek (The Ministry of Higher Education, Science, and Technology) of the Republic of Indonesia (through LPPM ITB, Indonesia) for partial financial support. HLP would like to thank the National Research and Innovation Agency (BRIN) for its financial support through the Postdoctoral Program. HLP and MLP would like to thank the members of the Theoretical Physics Groups at Institut Teknologi Bandung for their hospitality. MFARS is supported by the Second Century Fund (C2F), Chulalongkorn University, Thailand. We thank the anonymous reviewers for their helpful comments and suggestions.